The molarity of Cl- ion in an aqueous solution which was (w/V) 2% NaCl , 4% CaCl2 and 6% NH4Cl will be 
  1. 0.342
  2. 0.721
  3. 1.12
  4. 2.18

Grade:11

2 Answers

Sunil Kumar FP
9 years ago
w/v percentage means weight of solute in gm in 100ml of solute
2%NaCl has mole of Cl- =2/58.5
similarly 4% will have 2*4/111
6% nh4cl =6/53.5
total mole in 100 ml=2/58.5 +4/111 +6/53.5
in 1 lt of the solution=2.18 mole
molarity=2.18
Muhammad Anas
6 years ago
Here we knoww/v % means wt. of solute in g per 100ml of solute2%NaCl has mole of Cl- =1×(2/58.5)4% of CaCl2 will have Cl- = 2×(4/111)6% NH4CL has Cl- =1×(6/53.5)total mole in 100 ml=(2/58.5) +(8/111 )+(6/53.5)=0.218But in1000ml (1L) of the solution=(0.218/100)×1000=2.18molesMolarity=2.18mol/1L=2.18M